# Bootstrapping Autonomous Radars with Self-Supervised Learning

### v1

@@ 6.2 Ablating Augmentations
To better understand the value of each augmentation, we dissect the contribution of individual repurposed and radar-specific augmentations, as well as effect of removing the augmentations from the best combinations. Individual Augmentations. We start by comparing the effect of adding individual augmentations enumerated in sec. 4.4. We also list three other augmentations that we experimented with but did not yield beneficial results. They include two standard SSL vision augmentations: Cutout and Vertical Flip [18]. We also tested another radar-specific augmentation, Thresholding, whereby we created a 0-1 mask by setting a power (pixel-value) threshold for the radar heatmap.
@@
In our experiments, we pre-train Radical’s network using only one augmentation at a time, and fine-tune it with the 13K labeled dataset. The results are shown in Table 4. As a baseline for comparison, we also include the cross-modal only baseline in the table which does not use any augmentations. As seen, four out of the seven tested augmentations, namely rotation, RMM, center crop, and horizontal flip prove beneficial for the pre-training, while thresholding, cutout, and vertical flipping will on average reduce the performance across all three listed metrics. Few points; (1) Based on these results, we removed the three worst-performing augmentations from our final model. (2) we briefly discuss the reason behind each augmentation’s effectiveness. First, while the radar heatmaps are symmetric along the mid-point of the azimuth axis, they are certainly not so along the range axis. This is why the horizontal flip keeps the underlying structure of the radar data while the vertical flip fails to do so, and hurts the final performance as a result. Then, while threshold might seem an intuitive extension of similar quantization methods in vision, it fails to deliver a better performance in radar data due to radar data being already extremely sparse by nature. Finally, center cropping and rotation borrowed from vision seem to boost the network performance, we they still keep the underlying semantics of a radar heat map. Finally, RMM proves to be a useful MIMO radar-specific augmentation for promoting robustness to Doppler-induced phase noise in MIMO radars [26]. Augmentation Combinations. From the four useful augmentations above, we experimented with different combinations to find the best one. We conducted five experiments; in the first one we used all four augmentations, and in each of the subsequent experiments we removed one out of four augmentations. The results are shown in Table. 5. We note that all these combinations perform about equally well under AP, all close to Radical’s best results. On the other hand, under mAP and AP, the (RMM + Center Crop + Horizontal Flip) combination is the clear winner, and hence used in Radical’s final model. Parameters of the RMM Augmentation. Given the novelty of this augmentation, we explore various parameter configurations to identify the most effective combination. The results are summarized in Table 6. As a baseline, the first row shows the results without using RMM, while other rows show different combinations of hyper-parameters and following the notation introduced in sec. 4.4. As shown in the table, keeping antennas with probability and outperforms other cases across all three metrics. In other words, randomly omitting 10% of the antenna pairs in MIMO and introducing a modest phase noise with a range of leads to the version of RMM with the best results. Hence, RMM enhances radar pre-training best when it is applied in a non-aggressive manner that does not compromise the integrity of the radar data.
@@ 6.3 Qualitative Results
To demonstrate Radical’s performance in action, we show example qualitative results comparing our method to the Radatron baseline (random initialization) in Fig. 4. Throughout the figure, ground truth is shown as dotted green bounding boxes, while model predictions are in solid red. The first row shows scenes in camera front-view. The second row shows the ground truth in the BEV of the same scene overlaid on top of the radar heat map. There are a few points worth mentioning. First, quite a few of the failure cases, namely columns Fig. 4(i, ii, iii, v, and vi) in the baseline arise from scenarios where a car is detected, albeit its orientation and exact bounding box are missed. This is due to the low resolution and specular nature of radar. These failures are mostly rectified by Radical’s network as seen in the figure. Second, Radical performs better in scenarios where a car’s radar reflection might get occluded by other cars in the scene, as in Fig. 4(vii). Both these failure cases are well known in radar object detection systems, as shown by previous work [25, 39]. Radical overcomes these failures thanks to pre-training radars to learn radar priors like specularity and sparsity jointly with vision features, which additionally carry semantic information such as precise car location and orientations. Finally, we note that Radatron performs reasonably closer compared to Radical when detecting the approximate location of vehicles which is reflective of their relatively closer AP performance compared to mAP and AP. In other words, Radical’s strength, as noted in sec. 6.1, lies in its more precise box detection in complex situations, illustrated in Fig.4.
@@ 7 Conclusion
In this paper, we presented a self-supervised approach to radar object detection in the context of self-driving cars, harnessing the largely untapped potential of vast quantities of unlabeled radar data. Our extensive evaluations illustrate that Radical achieves superior performance over supervised baselines by effectively combining intra- and cross-modal self-supervised learning, and employing radar-specific as well as vision-inspired augmentation in the context of contrastive learning. It is our hope that these contributions are followed by future advancements in the field of automative radar.

### v3

@@ 6.2 Ablating augmentations
To better understand the value of each augmentation, we dissect the contribution of individual repurposed and radar-specific augmentations, as well as effect of removing the augmentations from the best combinations.
Individual augmentations. We first compare the effect of adding individual augmentations described in Sec. 4.4. We also list three augmentations that we experimented with but did not yield beneficial results. They include two standard SSL vision augmentations: Cutout and Vertical Flip [19]. We also tested another radar-specific augmentation, Thresholding, whereby we created a binary mask by setting a power (pixel-value) threshold for the radar heatmap.
@@
In our experiments, we pre-train the Radical net using one augmentation at a time, and fine-tune it with the 13K labeled dataset. The results are shown in Table 4. As a baseline for comparison, we also include the cross-modal only baseline in the table which does not use any augmentations. As seen, four out of the seven tested augmentations, namely rotation, RMM, center crop, and horizontal flip prove beneficial for pre-training. On the other hand, thresholding, cutout, and vertical flipping will on average be detrimental to performance across all three listed metrics. We make the following points. First, based on these results, we removed the three worst-performing augmentations from our final model. Second, we make the following observations regarding the effectiveness of each augmentation:
1.
While radar heatmaps are symmetric along the mid-point of the azimuth axis, they are certainly not so along the range axis. This is why horizontal flip retains the underlying structure of the radar data, while vertical flip fails to do so and hurts performance.
2.
While thresholding might seem an intuitive extension to similar quantization methods in vision, it fails to aid performance in radar data because radar data is already extremely sparse in nature.
3.
Center cropping and rotation borrowed from vision seem to boost Radical performance. They preserve the underlying semantics of radar heatmaps.
4.
RMM is a useful MIMO radar-specific augmentation.
Combining augmentations. Having found four individually useful augmentation in Table. 4, we next explore how to best combine them. To this end, we conduct five experiments; the first uses all four augmentations, while each subsequent experiments removes one out of four augmentations at a time. The results are shown in Table. 5. We note that all these combinations perform seemingly equally well under the AP50 metric. However, metrics mAP and AP75 reveal that the combination RMM + Center Crop + Horizontal Flip is the clear winner. We hence use it in Radical’s final model. Hyper-parameters of RMM augmentation. Having just introduced RMM augmentation in this work, we next explore the configuration space of its hyper-parameters in order to identify an initial performant recipe. Table 6 sweeps and (cf. Sec. 4.4). To establish a comparative baseline, the first row of Table 6 shows our three AP metrics without using RMM. We observe that keeping virtual antennas with probability and noise randomisation with results in best performance across the three metrics. This amounts to randomly omitting 10% of the antennas. While sizeable, we view this antenna masking as non-aggressive and preserving of the integrity of the radar data.
@@ 6.3 Qualitative results
We next present qualitative results and compare Radical to the supervised baseline Radatron. Fig. 4 shows groundtruth as dotted green bounding boxes, and model predictions in solid red. Fig. 4 consists of three rows: (1) upper row depicts front-view camera images, (2) middle row depicts Radatron’s bounding box predictions overlaid on top of groundtruth in BEV, and (3) bottom row depicts Radical’s bounding box predictions and groundtruth. We make the following observations. First, quite a few of the failure cases—namely columns i, ii, iii, v, and vi in Fig. 4—in the baseline arise from scenarios where a car is detected, albeit its orientation and exact bounding box are missed. This is due to the low-resolution and specular nature of radar. These failures are mostly rectified by Radical’s network as seen in the bottom row. Second, Radical performs better in scenarios where a car’s radar reflection might get occluded by other cars in the scene, as in Fig. 4(vii). Both these failure cases are well known in radar object detection systems, as shown by previous work [27, 43]. Radical overcomes these failures thanks to pre-training radars to learn radar priors like specularity and sparsity jointly with vision features, which additionally carry semantic information such as precise car location and orientations. Finally, we note that Radatron performs reasonably closer compared to Radical when detecting the approximate location of vehicles which is reflective of their relatively closer AP50 performance compared to mAP and AP75. In other words, Radical’s strength, as noted in Sec. 5.3, lies in its more precise box detection in complex situations, illustrated in Fig.4.
Controlled Fog Experiment. We evaluate Radical on fog data in Radatron dataset. Figure 5 shows Radical can detect cars accurately in fog and even outperform Radatron.
@@ 7 Conclusion
In this paper, we presented a self-supervised approach to radar object detection in the context of self-driving cars, harnessing the largely untapped potential of vast quantities of unlabeled radar data. Our extensive evaluations illustrate that Radical achieves superior performance over supervised baselines by effectively combining intra- and cross-modal self-supervised learning, and employing radar-specific as well as vision-inspired augmentations in the context of contrastive learning. It is our hope that these contributions are followed by future advancements in the field of automotive radar.